OpenStudy (anonymous):

Find the slope of the line through each pair of points. 1: (1.5,-0.5) ; (-2/3 , 1/3) 2: (-0.5, -0.5) ; (-3,-4) 3: (0, -0.5) ; (7/5,10) Pleas help me :(

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}\] is what you need for all of them

OpenStudy (anonymous):

Use the formula, for some slope m: \[m=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

it is arithmetic from here on in

OpenStudy (anonymous):

I use that formuala to find the whole answer? :o

OpenStudy (anonymous):

\((1.5,-0.5) ; (-2/3 , 1/3)\) you will need \[\frac{\frac{1}{3}-(-0.5)}{-\frac{2}{3}-1.5}\] you math teacher must hate you to give you fractions and decimals in the same problem. lets work this one out

OpenStudy (anonymous):

first of all we cannot use both fractions and decimals. lets convert each decimal to a fraction so we can do the arithmetic

OpenStudy (anonymous):

Yeah she does -__- and okay lets do this!

OpenStudy (anonymous):

\(-0.5=-\frac{1}{2}\) and \(1.5=\frac{3}{2}\) so now we can continue

OpenStudy (anonymous):

Okayy

OpenStudy (anonymous):

\[\frac{\frac{1}{3}-(-0.5)}{-\frac{2}{3}-1.5}\] \[\frac{\frac{1}{3}+\frac{1}{2}}{-\frac{2}{3}-\frac{3}{2}}\] \[=\frac{\frac{5}{6}}{-\frac{13}{6}}\] \[=-\frac{5}{6}\times \frac{6}{13}\] \[=-\frac{5}{13}\]

OpenStudy (anonymous):

hope all the steps are clear

OpenStudy (anonymous):

so that would be the slope? like the final answer?

OpenStudy (anonymous):

yes, that would be the slope unless i messed up the arithmetic

OpenStudy (anonymous):

i don't have to plug that into any other formula? (I'm asking cause my teacher does some many steps i dont understand)

OpenStudy (anonymous):

that wasn't enough work for you? there is only one formula for the slope, the one we both wrote above lets check the answer http://www.wolframalpha.com/input/?i=slope++%283%2F2%2C-1%2F2%29++%28-2%2F3+%2C+1%2F3%29

OpenStudy (anonymous):

Ohhhhh okay I got it thanks a bunch!!

OpenStudy (anonymous):

yw